OpenStudy (anonymous):

Caroline runs 3 1/2 miles in 22 1/4 minutes. How many miles does she run per minute? A. 3/22 B. 62/221 C. 14/89 D. 4/39

OpenStudy (anonymous):

First you have to get common denominators, and 1/2=2/4 so now its 3 2/4 and 22 1/4

OpenStudy (anonymous):

then you do 3 2/4 divided by 22 1/4

OpenStudy (anonymous):

Distance, Rate, Time. Use the distance formula. d =rt, d/r = t, d/t = r. You have the distance (3 1/2) and you have the time (22 1/4). You need to figure what the rate is. d/t = r

OpenStudy (anonymous):

DO YOUR OWN HOMEWORK.